# Benchmark stars for mean stellar density and surface gravity estimates of solar-type stars

### v1

@@ Abstract
Adding an independent estimate of the mean stellar density, $\rho_{\star}$, as a constraint in the analysis of stars that host transiting exoplanets can significantly improve the precision of the planet radius estimate in cases where the light curve is too noisy to yield an accurate value of the transit impact parameter, e.g. the light curves of Earth-size planets orbiting in the habitable zone of Sun-like stars that will be obtained by the PLATO mission. I have compiled a sample of 36 solar-type stars for which analysis of high-quality light curves together with constraints on the orbital eccentricity yield mean stellar density measurements with a median error of 2.3 per cent. Of these, 8 are in transiting exoplanet systems and 28 in eclipsing binary systems with very low mass companions that contribute $<0.1$ per cent of the total flux in the V band. A re-calibrated empirical relation for stellar mass as a function of Teff, $\rho_{\star}$ and [Fe/H] has been used to find mass estimates with a typical precision of 5.2 per cent for the stars in this sample. Examples are given of how this sample can be used to test the accuracy and precision of $\rho_{\star}$ and $\log g$ estimates from catalogues of stellar parameters for solar-type stars.
@@ 1 Introduction
[…]
@@ 2 Re-calibration of empirical mass and radius polynomial functions
Torres et al. (2010) derived empirical relations for the mass and radius of stars expressed as polynomial functions of their effective temperature, surface gravity and metallicity based on their own compilation of mass and radius measurements accurate to 3 per cent or better for stars in binary systems. Enoch et al. (2010) derived an empirical relation for the mass and radius of stars as polynomial functions of their mean stellar density, effective temperature and metallicity based on the same data. There are now over 350 eclipsing binary star systems with mass and radius measurements of comparable quality listed on the DEBCat website (Southworth, 2015). [^2] https://www.astro.keele.ac.uk/jkt/debcat/ The quantity and quality of measurements for late-type stars in long-period eclipsing binary systems has improved substantially compared to the data available to Torres et al. and Enoch et al. so I have used the data from DEBCat downloaded on 2025-03-05 to derive new empirical relations that are suitable for estimating the mass of a typical planet host star, i.e. a slowly-rotating, late-type main-sequence or subgiant star. I removed systems containing pre-main-sequence stars and stars outside the effective temperature range $3940\,{\rm K}<{\rm T}_{\rm eff}<7220\,{\rm K}$. This effective temperature range corresponds to a spectral type range F0 V to K9 V. This covers the full range of ${\rm T}_{\rm eff}$ for the benchmark stars discussed below and the majority of known planet host stars. I also removed metal-poor stars from the sample ([Fe/H] $<-1$) because the empirical relations defined by Torres et al. and Enoch et al. include only a linear term in [Fe/H] and the primary use for these empirical relations will be to estimate the properties of stars with solar-like metallicity. I computed the oblateness of the stars in each binary system using equation (43) from Chandrasekhar (1933) and excluded stars from the sample with an oblateness $>0.0025$ and all stars in systems with orbital periods $P<5$ days. This excludes stars that are significantly distorted by the gravity of their companions and stars that rotate rapidly because the tidal interaction between the stars is strong enough to force them to rotate synchronously with the orbit.
@@
For the dwarf components of the eclipsing red-giant + dwarf-star systems KIC 8410637 and KIC 5640750 I use the values of $T_{\rm eff,2}$ from Table 4 of Themeßl et al. (2018) rather than the values quoted in DEBCat. These latter values were taken from Table 3 of Themeßl et al. but are not accurate because the central surface brightness ratio was used instead of the average surface brightness ratio to estimate $T_{\rm eff,2}$ (Southworth, priv. comm.). [^3] The central surface brightness ratio is not the same as the average surface brightness ratio for stars with different limb-darkening profiles. For KIC 6131659, I use the masses and radii derived from my own analysis of the Kepler light curve described in Appendix C. The binary systems EBLM J0113+31 and EBLM J0608--59 have been excluded from the calibration data because these two systems have directly-measured mass ratios and so can be used as a test of the reliability of the method described below to estimate stellar mass and mean stellar density for the solar-type star in EBLM systems.
@@
These equations can be used to estimate the mass and radius of stars with Teff and $\log g$ within the range covered by the calibration sample, as shown in Fig. 1. Most of the stars in the calibration sample have metallicities in the range $-0.4<{\rm[Fe/H]}<+0.3$, so it is not advisable to use the relations to estimate the mass or radius of stars outside this range.
The root mean square (rms) of the residuals for equations (4) and (5) are 0.021 for $\log_{10}(M/M_{\odot})$ (5.0 per cent) and 0.011 for $\log_{10}(R/R_{\odot})$ (2.6 per cent), respectively. The rms of the residuals for equations (6) and (7) are 0.020 (4.8 per cent) and 0.008 (1.9 per cent), respectively. For comparison, the rms of the residuals using the original coefficients from Enoch et al. (2010) in equations (6) and (7) are 0.030 (7.2 per cent) and 0.014 (3.2 per cent), respectively. Equation (6) is the only one of these polynomial functions used in this study. The other polynomial functions are provided here for the convenience of other researchers interested in estimating the properties of FGK-type stars on or near the main sequence that have metallicities in the range $-0.4<{\rm[Fe/H]}<+0.3$.
@@ 3 Mean stellar density benchmark stars
CoRoT--1a 36352297 $6343\pm 74$ $1.197\pm 0.026$ $4.294\pm 0.003$ $0.04\pm 0.03$ $-0.177\pm 0.002$ 1
HAT--P--7 424865156 $6575\pm 34$ $1.513\pm 0.008$ $3.906\pm 0.001$ $0.28\pm 0.02$ $-0.709\pm 0.001$ 2
HD 189733a 256364928 $4875\pm 43$ $0.751\pm 0.007$ $4.665\pm 0.004$ $-0.08\pm 0.03$ $0.278^{+0.005}_{-0.004}$ 3
HD 209458a 420814525 $6126\pm 18$ $1.118\pm 0.005$ $4.331\pm 0.006$ $0.04\pm 0.01$ $-0.137^{+0.007}_{-0.008}$ 4
Kepler--1 399860444 $5875\pm 13$ $0.999\pm 0.004$ $4.466\pm 0.002$ $0.01\pm 0.01$ $0.041\pm 0.003$ 2
WASP--4a 402026209 $5496\pm 19$ $0.895\pm 0.004$ $4.513\pm 0.007$ $0.05\pm 0.01$ $0.088\pm 0.010$ 5
WASP--18 100100827 $6599\pm 48$ $1.343\pm 0.012$ $4.274\pm 0.003$ $0.22\pm 0.03$ $-0.183\pm 0.004$ 6
WASP--43 36734222 $4652\pm 201$ $0.689\pm 0.035$ $4.733\pm 0.010$ $-0.08\pm 0.08$ $0.362^{+0.010}_{-0.008}$ 7
KOI--205 271749758 $5237\pm 60$ $0.925\pm 0.033$ $4.576\pm 0.015$ $0.14\pm 0.12$ $0.190\pm 0.020$ 8
EBLM J0057--19 268529943 $5580\pm 150$ $1.067\pm 0.070$ $4.245\pm 0.015$ $0.23\pm 0.09$ $-0.275\pm 0.017$ $0.017$ 9
EBLM J0113+31 400048097 $6124\pm 50$ $1.029\pm 0.025$ $4.148\pm 0.006$ $-0.33\pm 0.15$ $-0.441\pm 0.007$ 10
EBLM J0123+38 186514178 $6182\pm 91$ $1.533\pm 0.086$ $3.809\pm 0.014$ $0.45\pm 0.07$ $-0.852\pm 0.016$ $0.022$ 9
EBLM J0228+05 422844353 $6912\pm 134$ $1.482\pm 0.093$ $4.103\pm 0.011$ $0.08\pm 0.08$ $-0.417\pm 0.010$ $0.011$ 11
EBLM J0500--46 161577376 $5788\pm 116$ $1.026\pm 0.058$ $4.180\pm 0.013$ $-0.15\pm 0.05$ $-0.381\pm 0.014$ $0.022$ 11
EBLM J0526--34 24397947 $6307\pm 128$ $1.332\pm 0.080$ $3.886\pm 0.010$ $-0.05\pm 0.07$ $-0.765\pm 0.008$ $0.025$ 11
EBLM J0540--17 46627823 $6290\pm 77$ $1.266\pm 0.067$ $4.003\pm 0.014$ $-0.04\pm 0.05$ $-0.601\pm 0.017$ $0.016$ 9
EBLM J0608--59 260128333 $6031\pm 56$ $1.098\pm 0.018$ $4.240\pm 0.010$ $0.01\pm 0.05$ $-0.322\pm 0.016$ 12
EBLM J0627--67 167201539 $6498\pm 117$ $1.217\pm 0.082$ $4.199\pm 0.011$ $-0.16\pm 0.13$ $-0.316\pm 0.007$ $0.027$ 11
EBLM J0719+25 458377744 $6026\pm 67$ $1.142\pm 0.059$ $4.198\pm 0.013$ $0.04\pm 0.05$ $-0.332\pm 0.015$ $0.017$ 9
EBLM J0941--31 25776767 $6504\pm 101$ $1.412\pm 0.081$ $3.942\pm 0.011$ $0.08\pm 0.07$ $-0.669\pm 0.010$ $0.017$ 9
EBLM J0955--39 45599777 $6340\pm 80$ $1.097\pm 0.061$ $4.400\pm 0.011$ $-0.24\pm 0.08$ $-0.037\pm 0.010$ $0.023$ 9
EBLM J1013+01 277712294 $5200\pm 80$ $0.900\pm 0.050$ $4.447\pm 0.009$ $0.09\pm 0.08$ $-0.010\pm 0.005$ $0.027$ 9
EBLM J1305--31 124095230 $5913\pm 64$ $1.192\pm 0.060$ $4.102\pm 0.010$ $0.20\pm 0.04$ $-0.466\pm 0.010$ $0.028$ 9
EBLM J1522+42 116380226 $5738\pm 64$ $1.043\pm 0.053$ $4.163\pm 0.013$ $-0.06\pm 0.04$ $-0.403\pm 0.017$ $0.023$ 9
EBLM J1928--38 469755925 $5687\pm 62$ $1.050\pm 0.053$ $4.147\pm 0.010$ $-0.01\pm 0.04$ $-0.426\pm 0.011$ $0.034$ 9
EBLM J1934--42 143291764 $5648\pm 68$ $1.061\pm 0.055$ $4.448\pm 0.012$ $0.29\pm 0.05$ $0.028\pm 0.014$ $0.022$ 9
EBLM J2040--41 291116020 $5790\pm 63$ $1.014\pm 0.052$ $4.168\pm 0.014$ $-0.21\pm 0.04$ $-0.402\pm 0.018$ $0.019$ 9
EBLM J2046+06 383303945 $6302\pm 70$ $1.199\pm 0.062$ $4.234\pm 0.010$ $0.00\pm 0.05$ $-0.266\pm 0.011$ $0.023$ 9
EBLM J2046--40 369456296 $5763\pm 75$ $1.223\pm 0.064$ $4.028\pm 0.009$ $0.34\pm 0.05$ $-0.572\pm 0.007$ $0.017$ 9
EBLM J2217--04 439837578 $5848\pm 85$ $1.221\pm 0.066$ $4.039\pm 0.009$ $0.27\pm 0.06$ $-0.556\pm 0.007$ $0.022$ 13
EBLM J2315+23 436520557 $6027\pm 66$ $1.175\pm 0.061$ $4.079\pm 0.009$ $0.02\pm 0.05$ $-0.504\pm 0.007$ $0.024$ 9
EBLM J2343+29 405532120 $4984\pm 87$ $0.844\pm 0.045$ $4.593\pm 0.010$ $0.11\pm 0.05$ $0.195\pm 0.009$ $0.020$ 9
EBLM J2359+44 177644756 $6799\pm 83$ $1.516\pm 0.081$ $3.978\pm 0.009$ $0.12\pm 0.05$ $-0.600\pm 0.006$ $0.020$ 9
EPIC 219654213 391137939 $6305\pm 110$ $1.239\pm 0.075$ $4.047\pm 0.009$ $-0.08\pm 0.09$ $-0.540\pm 0.005$ $0.019$ 14
NGTS--EB--7 238060327 $5770\pm 110$ $1.177\pm 0.076$ $4.085\pm 0.016$ $0.26\pm 0.12$ $-0.495\pm 0.019$ $0.011$ 15
WASP--30 9725627 $6190\pm 50$ $1.201\pm 0.065$ $4.193\pm 0.020$ $0.07\pm 0.08$ $-0.328\pm 0.028$ $0.005$ 16
@@ 3.1.3 HD 189733
Agol et al. (2010) have derived a precise value for the mean stellar density of HD 189733 assuming a circular orbit based on their analysis of six transits of HD 189733 b observed with Spitzer space telescope (Werner et al., 2004) at 8 $\mu$m ($\tilde{\rho}=1.894\pm 0.012\,\rho_{\odot}$). Agol et al. also observed six eclipses of HD 189733 b with the same instrument from which they derive a tight constraint on the value of $e\cos(\omega)=0.000050\pm 0.000094$. To account for the possibility of a slightly eccentric orbit, the radial velocity measurements from the following sources were analysed together with this constraint on $e\cos(\omega)$: 1 – Boisse et al. (2009, SOPHIE); 2 – Winn et al. (2006, HIRES); 3 – Moutou et al. (2020, SPIRou). Radial velocity measurements observed during the transit were excluded from the analysis. The values of $T_{0}$ and $P$ were fixed at the values measured by Krenn et al. (2023). The details of this analysis are otherwise very similar to those described for CoRoT-1 in Section 3.1.1. The results are given in Table 7. The median value of $\psi$ with the 15.9 per cent and 84.1 per cent percentiles points of the PPD is given for the error estimate because these are significantly different from each other, i.e. the PPD for this parameter is significantly skewed. The mean stellar density with the error estimate computed in the same way is $\rho_{\star}=\tilde{\rho}\times\psi=1.897^{+0.023}_{-0.016}\,\rho_{\odot}$.

### v2

@@ Abstract
Adding an independent estimate of the mean stellar density, $\rho_{\star}$, as a constraint in the analysis of stars that host transiting exoplanets can significantly improve the precision of the planet radius estimate in cases where the light curve is too noisy to yield an accurate value of the transit impact parameter, e.g. the light curves of Earth-size planets orbiting in the habitable zone of Sun-like stars that will be obtained by the PLATO mission. I have compiled a sample of 36 solar-type stars for which analysis of high-quality light curves together with constraints on the orbital eccentricity yield mean stellar density measurements with a median error of 2.3 per cent. Of these, 8 are in transiting exoplanet systems and 28 in eclipsing binary systems with very low mass companions that contribute <0.1<0.1 per cent of the total flux in the V band. A re-calibrated empirical relation for stellar mass as a function of Teff, $\rho_{\star}$ and [Fe/H] has been used to find mass estimates with a typical precision of 5.2 per cent for the stars in this sample. Examples are given of how this sample can be used to test the accuracy and precision of $\rho_{\star}$ and $\log g$ estimates from catalogues of stellar parameters for solar-type stars.
@@ 1 Introduction
$\mbox{$\rho_{\star}$}=\frac{3\mbox{M${}_{\star}$}}{4\pi R_{\star}^{3}}=\frac{3\pi}{GP^{2}(1+q)}\left(\frac{a}{\mbox{R${}_{\star}$}}\right)^{3},$ (1)
@@ 2 Re-calibration of empirical mass and radius polynomial functions
Torres et al. (2010) derived empirical relations for the mass and radius of stars expressed as polynomial functions of their effective temperature, surface gravity and metallicity based on their own compilation of mass and radius measurements accurate to 3 per cent or better for stars in binary systems. Enoch et al. (2010) derived an empirical relation for the mass and radius of stars as polynomial functions of their mean stellar density, effective temperature and metallicity based on the same data. There are now over 350 eclipsing binary star systems with mass and radius measurements of comparable quality listed on the DEBCat website (Southworth, 2015). [^2] https://www.astro.keele.ac.uk/jkt/debcat/ The quantity and quality of measurements for late-type stars in long-period eclipsing binary systems has improved substantially compared to the data available to Torres et al. and Enoch et al. so I have used the data from DEBCat downloaded on 2025-03-05 to derive new empirical relations that are suitable for estimating the mass of a typical planet host star, i.e. a slowly-rotating, late-type main-sequence or subgiant star. I removed systems containing pre-main-sequence stars and stars outside the effective temperature range $3940\,{\rm K}<{\rm T}_{\rm eff}<7220\,{\rm K}$. This effective temperature range corresponds to a spectral type range F0 V to K9 V. This covers the full range of ${\rm T}_{\rm eff}$ for the benchmark stars discussed below and the majority of known planet host stars. I also removed metal-poor stars from the sample ([Fe/H] <1<-1) because the empirical relations defined by Torres et al. and Enoch et al. include only a linear term in [Fe/H] and the primary use for these empirical relations will be to estimate the properties of stars with solar-like metallicity. I computed the oblateness of the stars in each binary system using equation (43) from Chandrasekhar (1933) and excluded stars from the sample with an oblateness >0.0025>0.0025 and all stars in systems with orbital periods $P<5$ days. This excludes stars that are significantly distorted by the gravity of their companions and stars that rotate rapidly because the tidal interaction between the stars is strong enough to force them to rotate synchronously with the orbit.
@@
For the dwarf components of the eclipsing red-giant + dwarf-star systems KIC 8410637 and KIC 5640750 I use the values of $T_{\rm eff,2}$ from Table 4 of Themeßl et al. (2018) rather than the values quoted in DEBCat. These latter values were taken from Table 3 of Themeßl et al. but are not accurate because the central surface brightness ratio was used instead of the average surface brightness ratio to estimate $T_{\rm eff,2}$ (Southworth, priv. comm.). [^3] The central surface brightness ratio is not the same as the average surface brightness ratio for stars with different limb-darkening profiles. For KIC 6131659, I use the masses and radii derived from my own analysis of the Kepler light curve described in Appendix C. The binary systems EBLM J0113+31 and EBLM J0608-59 have been excluded from the calibration data because these two systems have directly-measured mass ratios and so can be used as a test of the reliability of the method described below to estimate stellar mass and mean stellar density for the solar-type star in EBLM systems.
@@
These equations can be used to estimate the mass and radius of stars with Teff and $\log g$ within the range covered by the calibration sample, as shown in Fig. 1. Most of the stars in the calibration sample have metallicities in the range $0.4<{\rm[Fe/H]}<+0.3$, so it is not advisable to use the relations to estimate the mass or radius of stars outside this range.
The root mean square (rms) of the residuals for equations (4) and (5) are 0.021 for $\log_{10}(M/M_{\odot})$ (5.0 per cent) and 0.011 for $\log_{10}(R/R_{\odot})$ (2.6 per cent), respectively. The rms of the residuals for equations (6) and (7) are 0.020 (4.8 per cent) and 0.008 (1.9 per cent), respectively. For comparison, the rms of the residuals using the original coefficients from Enoch et al. (2010) in equations (6) and (7) are 0.030 (7.2 per cent) and 0.014 (3.2 per cent), respectively. Equation (6) is the only one of these polynomial functions used in this study. The other polynomial functions are provided here for the convenience of other researchers interested in estimating the properties of FGK-type stars on or near the main sequence that have metallicities in the range $0.4<{\rm[Fe/H]}<+0.3$.
@@ 3 Mean stellar density benchmark stars
CoRoT-1a 36352297 $6343\pm 74$ $1.209\pm 0.034$ $4.348\pm 0.004$ $0.04\pm 0.03$ $0.177\pm 0.002$ 1
HAT-P-7 424865156 $6575\pm 34$ $1.541\pm 0.034$ $4.028\pm 0.003$ $0.28\pm 0.02$ $0.709\pm 0.001$ 2
HD 189733a 256364928 $4875\pm 43$ $0.775\pm 0.018$ $4.587\pm 0.004$ $0.08\pm 0.03$ $0.278^{+0.005}_{-0.004}$ 3
HD 209458a 420814525 $6126\pm 18$ $1.136\pm 0.023$ $4.365\pm 0.006$ $0.04\pm 0.01$ $0.137^{+0.007}_{-0.008}$ 4
Kepler-1 399860444 $5875\pm 13$ $1.035\pm 0.020$ $4.471\pm 0.004$ $01\pm 0.01 0.0$$041\pm 0.003$ 2
WASP-4a 402026209 $5496\pm 19$ $0.947\pm 0.019$ $4.489\pm 0.007$ $05\pm 0.01 0.0$$088\pm 0.010$ 5
WASP-18 100100827 $6599\pm 48$ $1.362\pm 0.033$ $4.361\pm 0.004$ $0.22\pm 0.03$ $0.183\pm 0.004$ 6
WASP-43 36734222 $4652\pm 201$ $0.734\pm 0.040$ $4.635\pm 0.010$ $0.08\pm 0.08$ $0.362^{+0.010}_{-0.008}$ 7
KOI-205 271749758 $5237\pm 60$ $0.905\pm 0.038$ $4.550\pm 0.015$ $0.14\pm 0.12$ $0.190\pm 0.020$ 8
EBLM J0057-19 268529943 $5580\pm 150$ $1.067\pm 0.070$ $4.263\pm 0.014$ $0.23\pm 0.09$ $275\pm 0.017 0.0$170.017 9
EBLM J0113+31 400048097 $6124\pm 50$ $1.029\pm 0.025$ $4.148\pm 0.006$ $0.33\pm 0.15$ $0.441\pm 0.007$ 10
EBLM J0123+38 186514178 $6182\pm 91$ $1.533\pm 0.086$ $3.932\pm 0.012$ $0.45\pm 0.07$ $852\pm 0.016 0.0$220.022 9
EBLM J0228+05 422844353 $6912\pm 134$ $1.482\pm 0.093$ $4.216\pm 0.010$ $0.08\pm 0.08$ $417\pm 0.010 0.0$110.011 11
EBLM J0500-46 161577376 $5788\pm 116$ $1.026\pm 0.058$ $4.187\pm 0.011$ $0.15\pm 0.05$ $381\pm 0.014 0.0$220.022 11
EBLM J0526-34 24397947 $6307\pm 128$ $1.332\pm 0.080$ $3.969\pm 0.009$ $0.05\pm 0.07$ $765\pm 0.008 0.0$250.025 11
EBLM J0540-17 46627823 $6290\pm 77$ $1.266\pm 0.067$ $4.072\pm 0.012$ $0.04\pm 0.05$ $\pm 0.017 0.0160.016$ 9
EBLM J0608-59 260128333 $6031\pm 56$ $1.098\pm 0.018$ $4.240\pm 0.010$ $0.01\pm 0.05$ $0.322\pm 0.016$ 12
EBLM J0627-67 167201539 $6498\pm 117$ $1.217\pm 0.082$ $4.255\pm 0.010$ $0.16\pm 0.13$ $316\pm 0.007 0.0$270.027 11
EBLM J0719+25 458377744 $6026\pm 67$ $1.142\pm 0.059$ $4.236\pm 0.011$ $0.04\pm 0.05$ $332\pm 0.015 0.0$170.017 9
EBLM J0941-31 25776767 $6504\pm 101$ $1.412\pm 0.081$ $4.042\pm 0.009$ $0.08\pm 0.07$ $669\pm 0.010 0.0$170.017 9
EBLM J0955-39 45599777 $6340\pm 80$ $1.097\pm 0.061$ $4.427\pm 0.009$ $24\pm 0.08 0.0$$37\pm 0.010 0.0$230.023 9
EBLM J1013+01 277712294 $5200\pm 80$ $0.900\pm 0.050$ $4.416\pm 0.007$ $09\pm 0.08 0.0$$10\pm 0.005 0.0$270.027 9
EBLM J1305-31 124095230 $5913\pm 64$ $1.192\pm 0.060$ $4.153\pm 0.008$ $0.20\pm 0.04$ $466\pm 0.010 0.0$280.028 9
EBLM J1522+42 116380226 $5738\pm 64$ $1.043\pm 0.053$ $4.175\pm 0.012$ $0.06\pm 0.04$ $403\pm 0.017 0.0$230.023 9
EBLM J1928-38 469755925 $5687\pm 62$ $1.050\pm 0.053$ $4.161\pm 0.008$ $0.01\pm 0.04$ $426\pm 0.011 0.0$340.034 9
EBLM J1934-42 143291764 $5648\pm 68$ $1.061\pm 0.055$ $4.465\pm 0.011$ $29\pm 0.05 0.0$$8\pm 0.014 0.022$0.022 9
EBLM J2040-41 291116020 $5790\pm 63$ $1.014\pm 0.052$ $4.172\pm 0.013$ $0.21\pm 0.04$ $402\pm 0.018 0.0$190.019 9
EBLM J2046+06 383303945 $6302\pm 70$ $1.199\pm 0.062$ $4.287\pm 0.009$ $0.00\pm 0.05$ $266\pm 0.011 0.0$230.023 9
EBLM J2046-40 369456296 $5763\pm 75$ $1.223\pm 0.064$ $4.086\pm 0.007$ $0.34\pm 0.05$ $572\pm 0.007 0.0$170.017 9
EBLM J2217-04 439837578 $5848\pm 85$ $1.221\pm 0.066$ $4.096\pm 0.007$ $0.27\pm 0.06$ $556\pm 0.007 0.0$220.022 13
EBLM J2315+23 436520557 $6027\pm 66$ $1.175\pm 0.061$ $4.125\pm 0.007$ $0.02\pm 0.05$ $504\pm 0.007 0.0$240.024 9
EBLM J2343+29 405532120 $4984\pm 87$ $0.844\pm 0.045$ $4.544\pm 0.008$ $0.11\pm 0.05$ $195\pm 0.009 0.0$200.020 9
EBLM J2359+44 177644756 $6799\pm 83$ $1.516\pm 0.081$ $4.098\pm 0.006$ $0.12\pm 0.05$ $600\pm 0.006 0.0$200.020 9
EPIC 219654213 391137939 $6305\pm 110$ $1.239\pm 0.075$ $4.109\pm 0.007$ $0.08\pm 0.09$ $540\pm 0.005 0.0$190.019 14
NGTS-EB-7 238060327 $5770\pm 110$ $1.177\pm 0.076$ $4.131\pm 0.015$ $0.26\pm 0.12$ $495\pm 0.019 0.0$110.011 15
WASP-30 9725627 $6190\pm 50$ $1.201\pm 0.065$ $4.245\pm 0.019$ $0.07\pm 0.08$ $328\pm 0.028 0.0$050.005 16
@@ 3.1.3 HD 189733
[…]